### v1

# A generalization of Barannikov-Kontsevich theorem

Abstract
We study the twisted de Rham complex associated with a holomorphic function on a Kähler manifold whose critical point set is compact. We prove the $E_{1}$-degeneration of the Hodge-to-de Rham spectral sequence. It is a generalization of Barannikov-Kontsevich Theorem.
MSC:
Keywords: twisted de Rham complex, integrable mixed twistor $D$-modules, strictness, $E_{1}$-degeneration
1 Introduction
1.1 Twisted de Rham complexes
Let $X$ be a complex manifold. Let $f$ be a holomorphic function on $X$ such that the set of critical points $\mathop{\rm Cr}\nolimits(f)$ is compact.
Let $\Omega_{X}^{k}$ denote the sheaf of holomorphic $k$-forms on $X$. There exists the exterior derivative $d:\Omega^{k}_{X}\to\Omega^{k+1}_{X}$. The exterior product of $df$ induces morphisms $df:\Omega^{k}_{X}\to\Omega^{k+1}_{X}$, i.e., $\tau\longmapsto df\wedge\tau$.
Let $\lambda$ be a variable. Let $\Omega_{X}^{k}[\![\lambda]\!]$ denote the sheaf of formal power series with $\Omega_{X}^{k}$-coefficients. Namely, for any open subset $U$ of $X$, let $\Omega_{X}^{k}[\![\lambda]\!](U)$ denote the space of formal power series $\sum_{j\geq 0}\tau_{j}\lambda^{j}$, where $\tau_{j}\in\Omega_{X}^{k}(U)$. We obtain the differential $\lambda d+df:\Omega_{X}^{k}[\![\lambda]\!]\to\Omega_{X}^{k+1}[\![\lambda]\!]$ determined by $\tau\mapsto\lambda d\tau+df\wedge\tau$. It satisfies $(\lambda d+df)^{2}=0$. The complex $\Omega_{X}^{\bullet}[\![\lambda]\!]_{f}=(\Omega_{X}^{\bullet}[\![\lambda]\!],\lambda d+df)$ is called the twisted de Rham complex, or the formal twisted de Rham complex when we emphasize to consider the formal series. We obtain the cohomology group $H^{\ast}(X,\Omega^{\bullet}_{X}[\![\lambda]\!]_{f})$. It is naturally a ${\mathbb{C}}[\![\lambda]\!]$-module, where ${\mathbb{C}}[\![\lambda]\!]$ denotes the ring of formal power series with $\mathbb{C}$}-coefficients.
There are several different versions of twisted de Rham complex. For example, if $X$ and $f$ are algebraic, let $\Omega^{k}_{X^{\mathop{\rm alg}\nolimits}}$ denote the sheaf of algebraic $k$-forms on the algebraic variety $X^{\mathop{\rm alg}\nolimits}$ with Zariski topology, and it is also natural to consider the sheaf $\Omega_{X^{\mathop{\rm alg}\nolimits}}^{k}[\lambda]$ of polynomials with $\Omega_{X^{\mathop{\rm alg}\nolimits}}^{k}$-coefficients on $X^{\mathop{\rm alg}\nolimits}$. We obtain the algebraic twisted de Rham complex $\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}}[\lambda]_{f}=(\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}}[\lambda],\lambda d+df)$ and the cohomology group $H^{\ast}(X^{\mathop{\rm alg}\nolimits},\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}}[\lambda]_{f})$.
1.2 A basic question
By setting $\mathcal{F}_{\lambda}^{j}(\Omega^{\bullet}_{X}[\![\lambda]\!]_{f})=\lambda^{j}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ for any non-negative integer $j$, we obtain the filtered complex $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$. The associated complexes $\mathop{\rm Gr}\nolimits^{j}_{\mathcal{F}_{\lambda}}(\Omega^{\bullet}_{X}[\![\lambda]\!]_{f})=\mathcal{F}_{\lambda}^{j}/\mathcal{F}_{\lambda}^{j+1}$ $(j\geq 0)$ are isomorphic to the complex $(\Omega^{\bullet}_{X},df)$. Note that the cohomological support of $(\Omega^{\bullet}_{X},df)$ is contained in $\mathop{\rm Cr}\nolimits(f)$. Because $\mathop{\rm Cr}\nolimits(f)$ is compact, $H^{\ast}(X,(\Omega^{\bullet}_{X},df))$ is a finite dimensional complex vector space. There is the spectral sequence associated with the filtered complex, for which
$E_{1}^{p,q}=H^{p+q}\bigl{(}X,\mathop{\rm Gr}\nolimits^{p}_{\mathcal{F}_{\lambda}}(\Omega^{\bullet}_{X}[\![\lambda]\!]_{f})\bigr{)}=H^{p+q}(X,(\Omega_{X}^{\bullet},df)).$
We study the following question.
Question 1.1
Is the spectral sequence for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ degenerates at the $E_{1}$-level?
We can rephrase the condition in several ways.
Lemma 1.2
The spectral sequence degenerates at the $E_{1}$-level if and only if the following equivalent conditions are satisfied.
  • For any $\ell$ and $j$, the exact sequences $0\longrightarrow\mathop{\rm Gr}\nolimits_{\mathcal{F}_{\lambda}}^{j}\longrightarrow\Omega_{X}^{\bullet}[\![\lambda]\!]_{f}/\mathcal{F}_{\lambda}^{j+1}\longrightarrow\Omega_{X}^{\bullet}[\![\lambda]\!]_{f}/\mathcal{F}_{\lambda}^{j}\longrightarrow 0$ induce the following exact sequences:
    $0\longrightarrow H^{\ell}\bigl{(}X,\mathop{\rm Gr}\nolimits_{\mathcal{F}_{\lambda}}^{j}\bigl{(}\Omega_{X}^{\bullet}[\![\lambda]\!]_{f}\bigr{)}\bigr{)}\longrightarrow H^{\ell}\bigl{(}X,\Omega_{X}^{\bullet}[\![\lambda]\!]_{f}/\mathcal{F}_{\lambda}^{j+1}\bigr{)}\longrightarrow\\ H^{\ell}\bigl{(}X,\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}/\mathcal{F}_{\lambda}^{j}\bigr{)}\longrightarrow 0.$
  • For any $\ell$, $H^{\ell}(X,\Omega^{\bullet}_{X}[\![\lambda]\!]_{f})$ are isomorphic to $H^{\ell}(X,(\Omega^{\bullet}_{X},df))\otimes{\mathbb{C}}[\![\lambda]\!]$ as ${\mathbb{C}}[\![\lambda]\!]$-modules. In particular, the formal twisted de Rham cohomology groups $H^{\ell}(X,\Omega^{\bullet}_{X}[\![\lambda]\!]_{f})$ are free ${\mathbb{C}}[\![\lambda]\!]$-modules of finite rank.
Remark 1.3
It is easy to see that the second condition implies the first. For the convenience of readers, we shall recall that the $E_{1}$-degeneration is equivalent to the first condition in §3.1.3. We shall later explain that the first condition implies the second condition in a generalized context. (See Corollary 3.10.)
1.2.1 Classical results related to the question
There are two classical results related to this question. One appeared in the classical Hodge theory. As Deligne observed, if $X$ is projective and $f=0$, the desired $E_{1}$-degeneration follows from the Hodge decomposition of $H^{\ast}(X,{\mathbb{C}})$. Deligne generalized it to more general algebraic varieties by using mixed Hodge theory. This $E_{1}$-degeneration and its generalizations are not only deep results in the Hodge theory, but also useful for various applications including some vanishing theorems. Deligne and Illusie gave an alternative proof of the $E_{1}$-degeneration using the reduction to the positive characteristic.
The other appeared in the singularity theory, in particular, the study of Brieskorn lattices. If $f$ has only one critical point, the $E_{1}$-degeneration holds because the $\ell$-th cohomology group of $(\Omega^{\bullet}_{X},df)$ is 0 unless $\ell$ equals $\dim X$. In this case, the $E_{1}$-degeneration is an important starting point of the deep theory of primitive forms of Kyoji Saito. (See [19] and [20] for more backgrounds.)
1.3 Barannikov-Kontsevich Theorems and variations
The modern study of twisted de Rham complexes was opened by the celebrated theorem of Barannikov and Kontsevich.
Theorem 1.4 (Barannikov-Kontsevich)
The $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ holds in the case where $X$ and $f$ are quasi-projective.
This is a fundamental theorem in the study of the holomorphic Landau-Ginzburg model of the mirror symmetry. For example, it is essential in the proof of smoothness of some moduli spaces associated with Landau-Ginzburg models. (See [6].)
The original proof of Barannikov and Kontsevich was given by a generalization of the method of Deligne and Illusie. Indeed, Barannikov and Kontsevich proved the following theorem for the algebraic version of the twisted de Rham complexes, which implies Theorem 1.4.
Theorem 1.5 (Barannikov-Kontsevich)
Suppose that $f:X\to{\mathbb{C}}$ is a projective morphism of algebraic varieties. Then,
$\dim H^{j}(X^{\mathop{\rm alg}\nolimits},(\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}},df))=\dim H^{j}(X^{\mathop{\rm alg}\nolimits},(\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}},d+df))$
holds for any $j$ and for any complex number $\lambda$.
Theorem 1.5 implies that $H^{j}(X^{\mathop{\rm alg}\nolimits},\Omega_{X^{\mathop{\rm alg}\nolimits}}^{\bullet}[\lambda]_{f})$ are free ${\mathbb{C}}[\lambda]$-modules, and that the $E_{1}$-degeneration of the spectral sequence for the filtered complex $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}}[\lambda]_{f}$. In the setting of Theorem 1.4, there exists a projective morphism of algebraic varieties $F:Y\to{\mathbb{C}}$ with an open embedding $\iota:X\to Y$ such that $f=F\circ\iota$. Under the assumption that $\mathop{\rm Cr}\nolimits(f)$ is compact, the set $\mathop{\rm Cr}\nolimits(F)$ of critical points of $F$ is decomposed as $\mathop{\rm Cr}\nolimits(F)=\mathop{\rm Cr}\nolimits(f)\sqcup\bigl{(}\mathop{\rm Cr}\nolimits(F)\cap(Y\setminus X)\bigr{)}$. Then, we obtain the $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ from the $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{Y^{\mathop{\rm alg}\nolimits}}[\lambda]_{f}$.
The theorem of Barannikov-Kontsevich for the algebraic twisted de Rham complexes (Theorem 1.5) has attracted many mathematicians because of its significance in the non-commutative Hodge theory (see [6]), and because the theorem and its generalization are deeply related with various fields of mathematics. Indeed, alternative proofs for Theorem 1.5 with different methods have been found by Sabbah [15] using Hodge modules and microlocalization, and by Ogus and Vologodsky [14] using their non-abelian Hodge correspondence in positive characteristic. Later, Arinkin, Căldăraru and Hablicsek [1] revisited it in their study of Deligne-Illusie method from the viewpoint of derived algebraic geometry. Sabbah also studied generalizations to the case where $f$ is not necessarily projective but cohomologically tame [17]. See [2] and [12] for a generalization to the Kontsevich complexes.
The $L^{2}$-analogue of the twisted de Rham complex associated with $(X,f)$ has been also studied in [3, 7]. In particular, Li and Wen [7] studied the case where $X$ has a complete Kähler metric with bounded curvature, and $f$ is strongly elliptic, which is a kind of non-degeneracy condition at infinity. They established an analogue of Theorem 1.5 in this context.
1.4 Main result
In this paper, we shall study a generalization of Theorem 1.4. It is an affirmative answer to a question asked by Kontsevich to the author.
Theorem 1.6
If $X$ is Kähler, the $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega_{X}^{\bullet}[\![\lambda]\!]_{f}$ holds.
Note that $X$ can be a small neighbourhood of $\mathop{\rm Cr}\nolimits(f)$, and that we do not need any assumption on the behaviour of $(X,f)$ at infinity. It is our purpose to show that a local assumption around $\mathop{\rm Cr}\nolimits(f)$ is enough for the $E_{1}$-degeneration for the formal twisted de Rham complex, though global assumptions are useful to obtain stronger consequences as in Theorem 1.5.
Acknowledgement
I would like to thank Maxim Kontsevich for what he has brought to mathematics. This study is motivated by one of his questions. Theorem 1.6 is one of three topics in the author’s talk in the conference “Mathematics on the Crossroad of Centuries” held in 2024 September. The others will be explained elsewhere. A preliminary version of this paper was prepared for the conference in Rikkyo University in 2023 January. I heartily thank the organizers for the opportunity of the talks.
I thank Claude Sabbah for helpful discussions and for his kindness.
I am partially supported by the Grant-in-Aid for Scientific Research (A) (No. 21H04429), the Grant-in-Aid for Scientific Research (A) (No. 22H00094), the Grant-in-Aid for Scientific Research (A) (No. 23H00083), and the Grant-in-Aid for Scientific Research (C) (No. 20K03609), Japan Society for the Promotion of Science. I am also partially supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
2 Integrable mixed twistor $\mathcal{D}$-modules
2.1 Mixed twistor $\mathcal{D}$-modules
The theory of twistor $\mathcal{D}$-modules has been developed in [16, 18] and [9, 10, 11] as a twistor version of Hodge modules [21, 22] inspired by the principle called Simpson’s Meta theorem [23].
2.1.1 $\mathcal{R}_{X}$-triples
For any complex manifold $X$, we set $\mathcal{X}={\mathbb{C}}_{\lambda}\times X$. Let $p_{\lambda}:\mathcal{X}\to X$ denote the projection. Let $\mathcal{D}_{\mathcal{X}}$ denote the sheaf of holomorphic linear differential operators on $\mathcal{X}$, and let $\Theta_{X}$ denote the tangent sheaf of $X$. We obtain the sheaf of subalgebras $\mathcal{R}_{X}\subset\mathcal{D}_{\mathcal{X}}$ generated by $\lambda\cdot(p_{\lambda}^{\ast}\Theta_{X})$ over $\mathcal{O}_{\mathcal{X}}$. If $X$ is an open subset in ${\mathbb{C}}^{n}$, we have $\mathcal{R}_{X}=\mathcal{O}_{\mathcal{X}}\langle\lambda\partial_{1},\ldots,\lambda\partial_{n}\rangle$.
We set ${\boldsymbol{S}}=\bigl{\{}\lambda\in{\mathbb{C}}\,\big{|}\,|\lambda|=1\bigr{\}}$. Let $\sigma:{\boldsymbol{S}}\to{\boldsymbol{S}}$ be defined by $\sigma(\lambda)=-\lambda$. Let $\mathfrak{D}\mathfrak{b}_{{\boldsymbol{S}}\times X/{\boldsymbol{S}}}$ denote the sheaf of distributions on ${\boldsymbol{S}}\times X$ which are continuous with respect to $\boldsymbol{S}$}. (See [16, §0.5].) It is naturally an $\mathcal{R}_{X|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{R}_{X|{\boldsymbol{S}}\times X})$-module by the action $(P_{1}\otimes\sigma^{-1}(P_{2}))\cdot\tau=P_{1}\overline{\sigma^{-1}(P_{2})}\tau$. A sesqui-linear pairing of $\mathcal{R}_{X}$-modules $\mathcal{M}^{\prime}$ and $\mathcal{M}^{\prime\prime}$ is a morphism of $\mathcal{R}_{X|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{R}_{X|{\boldsymbol{S}}\times X})$-modules $\mathcal{M}^{\prime}_{|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{M}^{\prime\prime}_{|{\boldsymbol{S}}\times X})\to\mathfrak{D}\mathfrak{b}_{{\boldsymbol{S}}\times X/{\boldsymbol{S}}}$. Such a tuple $(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$ is called an $\mathcal{R}_{X}$-triple.
For $\mathcal{R}_{X}$-triples $\mathcal{T}_{i}=(\mathcal{M}_{i}^{\prime},\mathcal{M}_{i}^{\prime\prime},C_{i})$ $(i=1,2)$, a morphism $\mathcal{T}_{1}\to\mathcal{T}_{2}$ is defined to be a pair of $\mathcal{R}_{X}$-homomorphisms $\varphi^{\prime}:\mathcal{M}_{2}^{\prime}\to\mathcal{M}_{1}^{\prime}$ and $\varphi^{\prime\prime}:\mathcal{M}_{1}^{\prime\prime}\to\mathcal{M}_{2}^{\prime\prime}$ such that $C_{1}\circ(\varphi^{\prime}\times\mathop{\rm id}\nolimits)=C_{2}\circ(\mathop{\rm id}\nolimits\times\varphi^{\prime\prime})$. The category of $\mathcal{R}_{X}$-triples is an abelian category.
For an increasing filtration $W$ of $\mathcal{T}$ in the category of $\mathcal{R}_{X}$-triples, we have the increasing filtrations $W(\mathcal{M}^{\prime})$ and $W(\mathcal{M}^{\prime\prime})$ such that $W_{j}(\mathcal{T})=\bigl{(}\mathcal{M}^{\prime}/W_{-j-1}(\mathcal{M}^{\prime}),W_{j}(\mathcal{M}^{\prime\prime}),C_{j}\bigr{)}$, where $C_{j}$ denote the induced sesqui-linear pairings.
2.1.2 Direct image of $\mathcal{R}_{X}$-triples
Let $F:X\longrightarrow Y$ be a morphism of complex manifolds. We set $\omega_{\mathcal{X}}:=\lambda^{-\dim X}p_{X}^{\ast}\omega_{X}$, where $\omega_{X}$ denotes the canonical line bundle of $X$. Similarly, we set $\omega_{\mathcal{Y}}:=\lambda^{-\dim Y}p_{Y}^{\ast}\omega_{Y}$. We set $\mathcal{R}_{Y\leftarrow X}:=\omega_{\mathcal{X}}\otimes_{F^{-1}(\mathcal{O}_{\mathcal{Y}})}F^{-1}(\mathcal{R}_{Y}\otimes\omega_{\mathcal{Y}}^{-1})$. For any $\mathcal{R}_{X}$-module $\mathcal{M}$, we obtain the following $\mathcal{R}_{Y}$-modules:
$F^{j}_{\dagger}(\mathcal{M}):=R^{j}(\mathop{\rm id}\nolimits_{{\mathbb{C}}}\times F)_{!}\bigl{(}\mathcal{R}_{Y\leftarrow X}\otimes^{L}_{\mathcal{R}_{X}}\mathcal{M}\bigr{)}\quad(j\in{\mathbb{Z}}).$
For any $\mathcal{R}_{X}$-triple $\mathcal{T}=(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$, we obtain the $\mathcal{R}_{X}$-triples $F_{\dagger}^{j}(\mathcal{T})=(F_{\dagger}^{-j}(\mathcal{M}^{\prime}),F_{\dagger}^{j}(\mathcal{M}^{\prime\prime}),F_{\dagger}C)$ $(j\in{\mathbb{Z}})$ on $Y$. (See [16, §1.4].) When $\mathcal{T}$ is equipped with an increasing filtration $W$, let $W_{j+\ell}(F_{\dagger}^{j}(\mathcal{T}))$ denote the image of $F_{\dagger}^{j}(W_{\ell}\mathcal{T})\to F_{\dagger}^{j}(\mathcal{T})$ by which we obtain the filtration $W$ on $F_{\dagger}^{j}(\mathcal{T})$.
2.1.3 Pure twistor $\mathcal{D}$-modules
A polarizable pure twistor $\mathcal{D}$-module of weight $w$ on a complex manifold $X$ is an $\mathcal{R}_{X}$-triple satisfying some conditions. (See [18, 10].) Let $\mathop{\rm MT}\nolimits(X,w)$ denote the category of polarizable pure twistor $\mathcal{D}_{X}$-modules of weight $w$.
Theorem 2.1 ([16, 9, 10])
Let $F:X\to Y$ be a projective morphism of complex manifolds. For any $\mathcal{T}\in\mathop{\rm MT}\nolimits(X,w)$, we have $F^{j}_{\dagger}(\mathcal{T})\in\mathop{\rm MT}\nolimits(Y,w+j)$.
There exists the full subcategory $\mathop{\rm MT}\nolimits_{\mathop{\rm reg}\nolimits}(X,w)\subset\mathop{\rm MT}\nolimits(X,w)$ of regular polarizable pure twistor $\mathcal{D}_{X}$-modules of weight $w$. (See [16, 9] for the regularity condition.) We obtain the following generalization of Theorem 2.1 in the regular case.
Theorem 2.2 ([13])
Let $F:X\to Y$ be a morphism of complex manifolds. Let $\mathcal{T}\in\mathop{\rm MT}\nolimits_{\mathop{\rm reg}\nolimits}(X,w)$. Suppose that $X$ is Kähler, and that the support of $\mathcal{T}$ is proper over $Y$. Then, we have $F^{j}_{\dagger}(\mathcal{T})\in\mathop{\rm MT}\nolimits_{\mathop{\rm reg}\nolimits}(Y,w+j)$.
2.1.4 Mixed twistor $\mathcal{D}$-modules
A mixed twistor $\mathcal{D}$-module on $X$ is a filtered $\mathcal{R}_{X}$-triple $(\mathcal{T},W)$ and satisfying some additional conditions. (See [11, §7].) Let $\mathop{\rm MTM}\nolimits(X)$ denote the category of mixed twistor $\mathcal{D}$-modules on $X$. The following theorem is fundamental.
Theorem 2.3 ([11])
Let $F:X\to Y$ be a projective morphism of complex manifolds. For $(\mathcal{T},W)\in\mathop{\rm MTM}\nolimits(X)$, we have $(F_{\dagger}^{j}(\mathcal{T}),W)\in\mathop{\rm MTM}\nolimits(Y)$ for any $j$.
For any $(\mathcal{T},W)\in\mathop{\rm MTM}\nolimits(X)$, we have $\mathop{\rm Gr}\nolimits^{W}_{w}(\mathcal{T})\in\mathop{\rm MT}\nolimits(X,w)$. There exists the full subcategory $\mathop{\rm MTM}\nolimits_{\mathop{\rm reg}\nolimits}(X)\subset\mathop{\rm MTM}\nolimits(X)$ of mixed twistor $\mathcal{D}_{X}$-modules $(\mathcal{T},W)$ such that $\mathop{\rm Gr}\nolimits^{W}_{w}(\mathcal{T})\in\mathop{\rm MT}\nolimits_{\mathop{\rm reg}\nolimits}(X,w)$. Theorem 2.3 is generalized as follows in the regular case.
Theorem 2.4 ([13])
Let $F:X\to Y$ be a morphism of complex manifolds. Let $\mathcal{T}\in\mathop{\rm MTM}\nolimits_{\mathop{\rm reg}\nolimits}(X)$. Suppose that $X$ is Kähler, and that the support of $\mathcal{T}$ is proper over $Y$. Then, we have $F^{j}_{\dagger}(\mathcal{T})\in\mathop{\rm MTM}\nolimits_{\mathop{\rm reg}\nolimits}(Y)$.
2.1.5 Pure and mixed twistor $\mathcal{D}$-modules on a point
Let $\mathop{\rm pt}\nolimits$ denote the set of one point. An $\mathcal{R}_{\mathop{\rm pt}\nolimits}$-module is an $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$-module. Let $\sigma:\mathbb{P}^{1}\to\mathbb{P}^{1}$ be the anti-holomorphic map defined by $\sigma(\lambda)=(-\overline{\lambda})^{-1}$. Let $\sigma:{\mathbb{C}}_{\lambda}^{\ast}\to{\mathbb{C}}_{\lambda}^{\ast}$ and $\sigma:\mathbb{P}^{1}\setminus\{0\}\to{\mathbb{C}}_{\lambda}$ denote the induced maps. Let $\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime}$ be locally free $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$-modules of finite rank with $\mathop{\rm rank}\nolimits\mathcal{M}^{\prime}=\mathop{\rm rank}\nolimits\mathcal{M}^{\prime\prime}$. Let
$C:\mathcal{M}^{\prime}_{|{\mathbb{C}}_{\lambda}^{\ast}}\otimes\sigma^{\ast}(\mathcal{M}^{\prime\prime}_{|{\mathbb{C}}_{\lambda}^{\ast}})\to\mathcal{O}_{{\mathbb{C}}_{\lambda}^{\ast}}$
be an $\mathcal{O}_{{\mathbb{C}}_{\lambda}^{\ast}}$-homomorphism which is perfect in the sense the induced morphism $\Psi_{C}:\sigma^{\ast}(\mathcal{M}^{\prime\prime})_{|{\mathbb{C}}_{\lambda}^{\ast}}\to(\mathcal{M}^{\prime})^{\lor}_{|{\mathbb{C}}_{\lambda}^{\ast}}$ is an isomorphism, where $(\mathcal{M}^{\prime})^{\lor}={\mathcal{H}om}_{\mathcal{O}_{{\mathbb{C}}_{\lambda}}}(\mathcal{M}^{\prime},\mathcal{O}_{{\mathbb{C}}_{\lambda}})$. Such a triple $(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$ is called a smooth $\mathcal{R}_{\mathop{\rm pt}\nolimits}$-triple. For any smooth $\mathcal{R}_{\mathop{\rm pt}\nolimits}$-triple $\mathcal{T}=(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$, we obtain the locally free $\mathcal{O}_{\mathbb{P}^{1}}$-module $\Upsilon(\mathcal{T})$ by gluing $(\mathcal{M}^{\prime})^{\lor}$ on ${\mathbb{C}}_{\lambda}$ and $\sigma^{\ast}(\mathcal{M}^{\prime\prime})$ on $\mathbb{P}^{1}\setminus\{0\}$ with $\Psi_{C}$.
A polarizable pure twistor $\mathcal{D}$-module on $\mathop{\rm pt}\nolimits$ is a smooth $\mathcal{R}_{\mathop{\rm pt}\nolimits}$-triple $\mathcal{T}$ such that $\Upsilon(\mathcal{T})$ is isomorphic to a direct sum of $\mathcal{O}_{\mathbb{P}^{1}}(w)$. A mixed twistor $\mathcal{D}$-module on $\mathop{\rm pt}\nolimits$ is a smooth $\mathcal{R}_{\mathop{\rm pt}\nolimits}$-triple $\mathcal{T}$ with a weight filtration $W$ such that $\mathop{\rm Gr}\nolimits^{W}_{w}(\mathcal{T})$ are pure of weight $w$.
Let $a_{X}:X\to\mathop{\rm pt}\nolimits$ denote the canonical morphism. We set $\widetilde{\Omega}^{k}_{\mathcal{X}/{\mathbb{C}}}:=\lambda^{-k}(p_{\lambda}^{\ast}\Omega_{X}^{k})$. We have the exterior derivative $d:\widetilde{\Omega}^{k}_{\mathcal{X}/{\mathbb{C}}}\to\widetilde{\Omega}^{k+1}_{\mathcal{X}/{\mathbb{C}}}$. For a coherent $\mathcal{R}_{X}$-module $\mathcal{M}$ with compact support, we obtain the complex of sheaves $\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}$ on $\mathcal{X}$. We have
$a_{X\dagger}^{\ell}(\mathcal{M})=R^{\dim X+\ell}(\mathop{\rm id}\nolimits\times a_{X})_{\ast}\Bigl{(}\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}$
as the $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$-module.
Corollary 2.5
Suppose that $\mathcal{M}$ underlies a regular mixed twistor $\mathcal{D}$-module on $X$ with compact support. We also assume that $X$ is Kähler. Then, $a_{X\dagger}^{\ell}(\mathcal{M})$ are locally free $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$-modules.
This is closely related with the $E_{1}$-degeneration property. We consider the subcomplexes $\mathcal{F}_{\lambda}^{j}(\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})=\lambda^{j}\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}$. Because $\lambda^{j}:a^{\ell}_{X\dagger}(\mathcal{M})\to a^{\ell}_{X\dagger}(\mathcal{M})$ are monomorphisms, the following is exact for any $\ell$ and $j$:
$0\longrightarrow R^{\dim X+\ell}(\mathop{\rm id}\nolimits\times a_{X})_{\ast}\mathcal{F}_{\lambda}^{j}(\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\mathcal{M})\longrightarrow R^{\dim X+\ell}(\mathop{\rm id}\nolimits\times a_{X})_{\ast}(\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\mathcal{M})\longrightarrow\\ R^{\dim X+\ell}(\mathop{\rm id}\nolimits\times a_{X})_{\ast}\bigl{(}(\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\mathcal{M})/\mathcal{F}_{\lambda}^{j}\bigr{)}\longrightarrow 0.$ (1)
Hence, the following is an epimorphism for any $j$ and $\ell$:
$R^{\dim X+\ell}(\mathop{\rm id}\nolimits\times a_{X})_{\ast}\bigl{(}(\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\mathcal{M})\big{/}\mathcal{F}_{\lambda}^{j+1}\bigr{)}\longrightarrow R^{\dim X+\ell}(\mathop{\rm id}\nolimits\times a_{X})_{\ast}\bigl{(}(\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\mathcal{M})/\mathcal{F}_{\lambda}^{j}\bigr{)}.$ (2)
This means the $E_{1}$-degeneration of the spectral sequence associated with the filtration $\mathcal{F}_{\lambda}$ on $\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\mathcal{M}$.
2.2 Integrable mixed twistor $\mathcal{D}$-modules
We set $\widetilde{\mathcal{R}}_{X}=\mathcal{R}_{X}\langle\lambda^{2}\partial_{\lambda}\rangle\subset\mathcal{D}_{\mathcal{X}}$. If $X$ is an open subset in ${\mathbb{C}}^{n}$, we have $\widetilde{\mathcal{R}}_{X}=\mathcal{O}_{\mathcal{X}}\langle\lambda\partial_{1},\ldots,\lambda\partial_{n},\lambda^{2}\partial_{\lambda}\rangle$. By the identification ${\boldsymbol{S}}=\{e^{\sqrt{-1}\theta}\}$, we obtain the vector field $\partial_{\theta}$ on $\boldsymbol{S}$}.
Let $\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime}$ be $\widetilde{\mathcal{R}}_{X}$-modules. Let $\mathcal{T}=(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$ be an $\mathcal{R}_{X}$-triple. For any section $m^{\prime}$ of $\mathcal{M}^{\prime}$ for $U\subset{\boldsymbol{S}}\times X$, we set $\partial_{\theta}m^{\prime}=\sqrt{-1}\lambda\partial_{\lambda}m^{\prime}=(\sqrt{-1}\lambda\partial_{\lambda}-\sqrt{-1}\overline{\lambda}\partial_{\overline{\lambda}})m^{\prime}$. Similarly, $\partial_{\theta}m^{\prime\prime}$ is defined for a section $m^{\prime\prime}$ of $\mathcal{M}^{\prime\prime}$. The $\mathcal{R}$-triple $\mathcal{T}$ is called integrable if
$\partial_{\theta}C(m^{\prime},\overline{\sigma^{-1}(m^{\prime\prime})})=C(\partial_{\theta}m^{\prime},\overline{\sigma^{-1}(m^{\prime\prime})})+C(m^{\prime},\overline{\sigma^{-1}(\partial_{\theta}m^{\prime\prime})}).$
(See [11, §2.1.5] for integrable $\mathcal{R}_{X}$-triples, which originally goes back to [16].) An integrable $\mathcal{R}_{X}$-triple is called $\widetilde{\mathcal{R}}_{X}$-triple. A morphism of $\widetilde{\mathcal{R}}_{X}$-triples $\mathcal{T}_{i}=(\mathcal{M}^{\prime}_{i},\mathcal{M}^{\prime\prime}_{i},C)$ $(i=1,2)$ is defined to be a morphism $(\varphi^{\prime},\varphi^{\prime\prime})$ of $\mathcal{R}_{X}$-triples such that $\varphi^{\prime}$ and $\varphi^{\prime\prime}$ are $\widetilde{\mathcal{R}}_{X}$-homomorphisms. For a morphism of complex manifolds $F:X\to Y$ and for an $\widetilde{\mathcal{R}}_{X}$-triple $\mathcal{T}$, the $\mathcal{R}_{Y}$-triples $F^{j}_{\dagger}(\mathcal{T})$ are naturally $\widetilde{\mathcal{R}}_{Y}$-triples.
An integrable mixed twistor $\mathcal{D}$-module on $X$ is a filtered $\widetilde{\mathcal{R}}_{X}$-triple $(\mathcal{T},W)$ satisfying some conditions. (See [11, §7.2.3].) Let $\mathop{\rm MTM}\nolimits_{\mathop{\rm reg}\nolimits}^{\mathop{\rm int}\nolimits}(X)$ denote the category of integrable mixed twistor $\mathcal{D}_{X}$-modules whose underlying mixed twistor $\mathcal{D}_{X}$-modules are regular. Let $\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$ denote the full subcategory of $\widetilde{\mathcal{R}}_{X}$-modules underlying regular integrable mixed twistor $\mathcal{D}_{X}$-modules, i.e., an $\widetilde{\mathcal{R}}_{X}$-module $\mathcal{M}^{\prime\prime}$ is an object of $\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$ if and only if there exists $((\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C),W)\in\mathop{\rm MTM}\nolimits^{\mathop{\rm int}\nolimits}_{\mathop{\rm reg}\nolimits}(X)$.
2.3 $\widetilde{\mathcal{R}}_{X}$-modules induced by Hodge modules
Let $\mathcal{D}_{X}$ denote the sheaf of holomorphic linear differential operators on $X$. Let $F_{j}(\mathcal{D}_{X})$ denote the subsheaf of differential operators of degree at most $j$. We set $R^{F}(\mathcal{D}_{X}):=\sum_{j\in{\mathbb{Z}}}\lambda^{j}F_{j}(\mathcal{D}_{X})$ and $\widetilde{R}^{F}(\mathcal{D}_{X}):=R^{F}(\mathcal{D}_{X})\langle\lambda^{2}\partial_{\lambda}\rangle$.
Let $M$ be a regular holonomic $\mathcal{D}_{X}$-module. Let $F(M)$ be a good filtration of $M$. We obtain $R^{F}(\mathcal{D}_{X})$-module $R^{F}(M)=\sum_{j\in{\mathbb{Z}}}\lambda^{j}F_{j}(M)$. It is naturally an $\widetilde{R}^{F}(\mathcal{D}_{X})$-module. By the analytification, it induces an $\widetilde{\mathcal{R}}_{X}$-module denoted by $\mathcal{R}^{F}(M)$. In this way, we obtain a functor from the category of good filtered regular holonomic $\mathcal{D}_{X}$-modules to the category of $\widetilde{\mathcal{R}}_{X}$-modules.
Lemma 2.6
If $(M,F)$ is a filtered regular holonomic $\mathcal{D}$-module underlying a mixed Hodge module, we have $\mathcal{R}^{F}(M)\in\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$. (See [21, 22] for Hodge modules.)
Proof  There exists a natural functor from the category of mixed Hodge modules on $X$ to $\mathop{\rm MTM}\nolimits^{\mathop{\rm int}\nolimits}_{\mathop{\rm reg}\nolimits}(X)$ as explained in [11, §13.5]. In the level of filtered $\mathcal{D}$-modules, it is given as above.
3 Main Theorem
3.1 Preliminary
Let $X$ be a complex manifold. Let $\mathcal{M}\in\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$ be an $\widetilde{\mathcal{R}}_{X}$-module underlying an integrable regular mixed twistor $\mathcal{D}$-module induced by a mixed Hodge module. Let $\mathop{\rm Ch}\nolimits(\Xi_{\mathop{\rm DR}\nolimits}\mathcal{M})\subset T^{\ast}X$ denote the characteristic variety of the underlying $\mathcal{D}_{X}$-module $\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M})$. Let $0_{X}:X\to T^{\ast}X$ denote the 0-section. We assume the following.
  • The set $\mathop{\rm Cr}\nolimits(f)$ is compact.
  • Any $c\neq 0$ is a regular value of $f$.
  • $\mathop{\rm Ch}\nolimits(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M}))\cap df(X)\subset 0_{X}(\mathop{\rm Cr}\nolimits(f))$.
The second condition implies $\mathop{\rm Cr}\nolimits(f)\subset f^{-1}(0)$. Because the characteristic varieties are cone, the third condition implies $\mathop{\rm Ch}\nolimits(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M}))\cap(\alpha df)(X)\subset 0_{X}(\mathop{\rm Cr}\nolimits(f))$ for any non-zero constant $\alpha$.
3.1.1 Cohomology group of the restriction to $\lambda=0$
Let $\mathcal{L}(f)$ denote the $\widetilde{\mathcal{R}}_{X}$-module given by $\mathcal{O}_{\mathcal{X}}$ with the meromorphic integrable connection $d+d(\lambda^{-1}f)$. We obtain the $\widetilde{\mathcal{R}}_{X}$-module $\mathcal{M}_{f}=\mathcal{M}\otimes\mathcal{L}(f)$ on $\mathcal{X}$, and the complex of sheaves $\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}$ on $\mathcal{X}$. We recall that $\mathcal{M}_{f}\in\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$. Let $\iota_{\lambda}:\{0\}\times X\to{\mathbb{C}}_{\lambda}\times X$ denote the inclusion. We obtain the complex of coherent $\mathop{\rm Sym}\nolimits\Theta_{X}$-modules $\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$. It induces a complex of coherent $\mathcal{O}_{T^{\ast}X}$-modules denoted by $\bigl{(}\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})\bigr{)}^{\sim}$.
Lemma 3.1
The cohomological support of $\bigl{(}\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})\bigr{)}^{\sim}$ is contained in $0_{X}(\mathop{\rm Cr}\nolimits(f))$. As a result,
$H^{\ast}\bigl{(}X,\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})\bigr{)}$
are finite dimensional.
Proof  Let $(\iota_{\lambda}^{\ast}\mathcal{M})^{\sim}$ and $(\iota_{\lambda}^{\ast}\mathcal{L}(f))^{\sim}$ denote the coherent $\mathcal{O}_{T^{\ast}X}$-modules induced by $\iota_{\lambda}^{\ast}\mathcal{M}$ and $\iota_{\lambda}^{\ast}\mathcal{L}(f)$, respectively. The support of $(\iota_{\lambda}^{\ast}\mathcal{M})^{\sim}$ is the characteristic variety $\mathop{\rm Ch}\nolimits(\Xi_{\mathop{\rm DR}\nolimits}\mathcal{M})$. The support of $(\iota_{\lambda}^{\ast}\mathcal{L}(f))^{\sim}$ is the image of $df(X)$. The support of $(\iota_{\lambda}^{\ast}\mathcal{M}_{f})^{\sim}$ is $df(X)+\mathop{\rm Ch}\nolimits(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M}))$ in $T^{\ast}X$.
Let $\omega_{X}$ denote the canonical bundle of $X$. Let $\mathop{\rm Sym}\nolimits\Theta_{X}\otimes\bigwedge^{\bullet}\Theta_{X}\otimes\omega_{X}$ be the Koszul resolution of $\mathcal{O}_{X}$ by $\mathop{\rm Sym}\nolimits\Theta_{X}$-free modules. It induces an $\mathcal{O}_{T^{\ast}X}$-free resolution $\mathcal{O}_{T^{\ast}X}\otimes\pi^{\ast}\bigl{(}\bigwedge^{\bullet}\Theta_{X}\otimes\omega_{X}\bigr{)}$ of $0_{X\ast}(\mathcal{O}_{X})$, where $\pi:T^{\ast}X\to X$ denotes the projection.
We have
$\bigl{(}\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})\bigr{)}^{\sim}\simeq(\iota_{\lambda}^{\ast}\mathcal{M}_{f})^{\sim}\otimes_{\mathcal{O}_{T^{\ast}X}}\Bigl{(}\mathcal{O}_{T^{\ast}X}\otimes\pi^{\ast}\Bigl{(}\bigwedge^{\bullet}\Theta_{X}\otimes\omega_{X}\Bigr{)}\Bigr{)}.$
Hence, the cohomological support is contained in the intersection of the support of $(\iota_{\lambda}^{\ast}\mathcal{M}_{f})^{\sim}$ and the $0_{X}(X)$, which is contained in $0_{X}(\mathop{\rm Cr}\nolimits(f))$.
3.1.2 Cohomology group of the vanishing cycle sheaf
Let $\iota_{f}:X\to X\times{\mathbb{C}}_{t}$ be the graph embedding, i.e., $\iota_{f}(x)=(x,f(x))$. There exists the $V$-filtration $V(\iota_{f\dagger}(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M})))$ along $t$. We obtain the regular holonomic $\mathcal{D}_{X}$-module
$\phi_{f}(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M})):=\bigoplus_{-1<a\leq 0}\mathop{\rm Gr}\nolimits^{V}_{a}\bigl{(}\iota_{f\dagger}\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M})\bigr{)}.$
Lemma 3.2
The support of $\phi_{f}(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M}))$ is contained in the compact subset $\mathop{\rm Cr}\nolimits(f)$. As a result, the cohomology group $H^{\ast}\bigl{(}X,\phi_{f}(\Xi_{\mathop{\rm DR}\nolimits}\mathcal{M})\otimes\Omega^{\bullet}_{X}\bigr{)}$ is finite dimensional.
Proof  The third condition implies that $\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M})$ is non-characteristic to the hypersurfaces $f^{-1}(c)$ on $X\setminus\mathop{\rm Cr}\nolimits(f)$. Hence, $\phi_{f}(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M}))=0$ on $X\setminus\mathop{\rm Cr}\nolimits(f)$.
3.1.3 $E_{1}$-degeneration and long exact sequences
We recall that the $E_{1}$-degeneration condition of a spectral sequence associated with a filtered complex for the convenience of readers. Let $\mathcal{A}$ be an abelian category. Let $(K^{\bullet},d)$ be a complex with a decreasing filtration $F^{\bullet}(K^{\bullet})$. To simplify the notation, we set $F^{p,p+r}(K^{j}):=F^{p}(K^{j})/F^{p+r}(K^{j})$.
Lemma 3.3
Let $r_{0}\in{\mathbb{Z}}_{\geq 1}$. The following conditions are equivalent.
$A(r_{0})$:
$H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $0\leq r\leq r_{0}$.
$B(r_{0})$:
$H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+r-1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $0\leq r\leq r_{0}$.
Proof  It is easy to see that $B(r_{0})$ implies $A(r_{0})$. Suppose that $A(r_{0})$ holds. We shall prove $H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+r-1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $0\leq r\leq r_{0}$, by an induction on $r$. We have the following commutative diagram:
[…]
By $A(r_{0})$, we have $b=0$. By the assumption of the induction on $r$, $c$ is a monomorphism. Hence, we obtain $a=0$, and the induction can proceed.
Recall that the spectral sequence $E_{r}^{p,q}$ for the filtered complex $F^{\bullet}(K^{\bullet})$ is given as
$E_{r}^{p,q}=Z_{r}^{p,q}\big{/}(Z_{r-1}^{p+1}+dZ^{p-r+1,q+r-2}_{r-1})$
by setting $Z_{r}^{p,q}=\mathop{\rm Ker}\nolimits\Bigl{(}d:F^{p}K^{p+q}\longrightarrow F^{p,p+r}(K^{p+q+1})\Bigr{)}$. (See [4] for more details.) There exist the natural morphisms $d_{r}:E_{r}^{p,q}\to E_{r}^{p+r,q-r+1}$ such that $d_{r}\circ d_{r}=0$, induced by $d:F^{p}K^{p+q}\to F^{p}K^{p+q+1}$. There exist natural isomorphisms
$E_{r+1}^{p,q}\simeq\mathop{\rm Ker}\nolimits\Bigl{(}E_{r}^{p,q}\to E_{r}^{p+r,q-r+1}\Bigr{)}\Big{/}\mathop{\rm Im}\nolimits\Bigl{(}E_{r}^{p-r,q+r-1}\to E_{r}^{p,q}\Bigr{)}.$
Lemma 3.4
We have $d_{r}=0$ for any $1\leq r\leq r_{0}-1$ if and only if $A(r_{0})$ holds.
Proof  We shall use an induction on $r_{0}$. Suppose that $d_{r}=0$ $(1\leq r\leq r_{0}-2)$ and that $A(r_{0}-1)$ holds. We have $E_{r}^{p,q}=E_{1}^{p,q}=H^{p+q}(F^{p,p+1}(K^{\bullet}))$ for $1\leq r\leq r_{0}-1$. We consider
[…]
By $A(r_{0}-1)$, $c$ is a monomorphism, and we have $b\circ a=0$. There exists a unique morphism $\varphi:H^{j}(F^{p,p+1})\to H^{j+1}(F^{p+r_{0}-1,p+r_{0}})$ such that $c\circ\varphi=b$. By the construction, $\varphi=d_{r_{0}-1}$. Hence, $A(r_{0})$ holds if and only if $d_{r_{0}-1}=0$.
We say that the spectral sequence degenerates at the $E_{1}$-level if $d_{r}=0$ for any $r\geq 1$. We obtain the following proposition.
Proposition 3.5
The $E_{1}$-degeneration holds if and only if one of the following equivalent conditions holds.
  • $H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+r-1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $1\leq r$.
3.2 Refinement of the Theorem 1.6
3.2.1 Main Theorem
Let us explain a refined statement of Theorem 1.6. We define the filtration $\mathcal{F}_{\lambda}^{k}(\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})=\lambda^{k}(\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ for $k\in{\mathbb{Z}}_{\geq 0}$. Note that $\mathop{\rm Gr}\nolimits_{\mathcal{F}_{\lambda}}^{k}\bigl{(}\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}$ is isomorphic to $\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ for any $k\geq 0$. Let us state the main theorem of this paper.
Theorem 3.6
The $E_{1}$-degeneration holds for the filtered complex $\mathcal{F}_{\lambda}(\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ with respect to the push-forward by the projection $\mathcal{X}\to{\mathbb{C}}_{\lambda}$. In other words, we obtain the following exact sequences for any $\ell$ and $j$:
$0\to H^{j}\Bigl{(}X,\mathop{\rm Gr}\nolimits^{\mathcal{F}_{\lambda}}_{\ell}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}\Bigr{)}\to H^{j}\Bigl{(}X,\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}\big{/}\mathcal{F}_{\lambda}^{\ell}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}\Bigr{)}\\ \to H^{j}\Bigl{(}X,\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}\big{/}\mathcal{F}_{\lambda}^{\ell+1}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}\Bigr{)}\to 0.$ (3)
We obtain Theorem 1.6 from Theorem 3.6 as the special case $\mathcal{M}=\mathcal{O}_{\mathcal{X}}$. We shall also prove the following.
Theorem 3.7
We have $\dim H^{j}\bigl{(}X,\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})\bigr{)}=\dim H^{j}\bigl{(}X,\phi_{f}(\Xi_{\mathop{\rm DR}\nolimits}\mathcal{M})\otimes\Omega^{\bullet}_{X}\bigr{)}$ for any $j$.
Corollary 3.8
In particular, we have $\dim H^{j}\bigl{(}X,(\Omega^{\bullet}_{X},df)\bigr{)}=\dim H^{j}(X,\Omega^{\bullet}_{X}\otimes\phi_{f}(\mathcal{O}_{X}))$.
Corollary 3.9
There exists an isomorphism
[…]
3.2.2 Completion
We naturally regard $\bigl{(}\mathcal{M}_{f}\bigl{/}\mathcal{F}^{j}_{\lambda}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}$ as the complexes of sheaves on $X$. We obtain the following complex of sheaves on $X$:
[…]
Corollary 3.10
There exists an isomorphism
$H^{j}\Bigl{(}X,\widehat{\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}}\Bigr{)}\simeq H^{j}\Bigl{(}X,\iota_{\lambda}^{\ast}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}\Bigr{)}\otimes{\mathbb{C}}[\![\lambda]\!].$
In particular, $H^{j}\Bigl{(}X,\widehat{\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}}\Bigr{)}$ are free ${\mathbb{C}}[\![\lambda]\!]$-modules of finite rank.
Proof  Let $\Omega^{0,q}_{X}$ denote the sheaf of smooth $(0,q)$-forms on $X$. We obtain the following double complex
[…]
Let $\mathop{\rm Tot}\nolimits\Bigl{(}\bigl{(}\widehat{\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}}\bigr{)}\widehat{\otimes}_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}\Bigr{)}$ denote the total complex.
Lemma 3.11
The natural morphism is a fine resolution:
$\widehat{\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}}\longrightarrow\mathop{\rm Tot}\nolimits\Bigl{(}\bigl{(}\widehat{\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}}\bigr{)}\widehat{\otimes}_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}\Bigr{)}$. (5)
Proof  It is enough to prove that (5) is a quasi-isomorphism locally around any point of $X$. Recall that the sheaf of $C^{\infty}$-functions on $X$ is flat over $\mathcal{O}_{X}$ according to [8].
Let $G$ be any pseudo-coherent $\mathcal{O}_{X}$-module. (See [5, Appendix A] for pseudo-coherent sheaves.) We have $G\otimes^{L}_{\mathcal{O}_{X}}\Omega^{0,q}_{X}\simeq G\otimes_{\mathcal{O}_{X}}\Omega^{0,q}_{X}$. Let $F_{\bullet}\to G$ be a free resolution of $G$. The natural morphisms $F_{\bullet}\to\mathop{\rm Tot}\nolimits(F_{\bullet}\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X})\to G\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}$ are quasi-isomorphisms. Hence, $G\to G\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}$ is a quasi-isomorphism.
Let $\mathcal{G}$ be a pseudo-coherent $\mathcal{O}_{\mathcal{X}}$-module flat over $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$. We naturally regard $\mathcal{G}/\lambda^{j}\mathcal{G}$ as $\mathcal{O}_{X}$-modules. Let $\pi_{j+1}:\mathcal{G}/\lambda^{j+1}\mathcal{G}\to\mathcal{G}/\lambda^{j}\mathcal{G}$ denote the projections. We obtain the quasi-isomorphisms
$\mathcal{G}/\lambda^{j}\mathcal{G}\to(\mathcal{G}/\lambda^{j}\mathcal{G})\otimes_{\mathcal{O}_{X}}\Omega_{X}^{0,\bullet}.$
For any open subset $U\subset X$, the morphisms $H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}\to H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}$ are surjective because $\iota_{\lambda}^{\ast}(\mathcal{G})\otimes\Omega^{0,q}_{X}$ is fine. Let $a_{j}^{q+1}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j}\mathcal{G})\otimes\Omega^{0,q+1}\bigr{)}$ $(j=1,2,\ldots)$ be sections such that $\overline{\partial}(a_{j}^{q+1})=0$ and $\pi_{j+1}(a_{j+1}^{q+1})=a_{j}^{q+1}$. Let us construct $b_{j}^{q}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j})\otimes\Omega^{0,q}_{X}\bigr{)}$ $(j=1,2,\ldots)$ such that $\overline{\partial}b_{j}^{q}=a_{j}^{q}$ and $\pi_{j+1}(b_{j+1}^{q})=b_{j}^{q}$ inductively on $j$. Suppose that we have already constructed $b_{j}^{q}$. There exists $c_{j+1}^{q}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}$ such that $\pi_{j+1}(c_{j+1}^{q})=b_{j}^{q}$. We obtain
$d_{j+1}^{q+1}=a_{j+1}^{q+1}-\overline{\partial}c_{j+1}^{q}\in H^{0}\bigl{(}U,(\lambda^{j}\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q+1}\bigr{)}\simeq H^{0}\bigl{(}U,\iota_{\lambda}^{\ast}\mathcal{G}\otimes\Omega^{0,q+1}\bigr{)}$
such that $\overline{\partial}(d_{j+1}^{q+1})=0$. There exists $e_{j+1}^{q}\in H^{0}\bigl{(}U,(\lambda^{j}\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}$ such that $\overline{\partial}e_{j+1}^{q}=d_{j+1}^{q+1}$. By setting $b_{j+1}^{q}=c_{j+1}^{q}+e_{j+1}^{q}$, the induction can proceed.
As a result, the natural morphism
[…]
is a quasi-isomorphism. Then, we obtain the claim of Lemma 3.11.
We obtain Corollary 3.10 from the following lemma.
Lemma 3.12
The natural morphisms
[…]
are isomorphisms.
Proof  Let $\pi_{\ell+1}$ denote the projection induced by $\mathcal{M}_{f}/\mathcal{F}_{\lambda}^{\ell+1}\mathcal{M}_{f}\to\mathcal{M}_{f}/\mathcal{F}_{\lambda}^{\ell}\mathcal{M}_{f}$. Let $d$ denote the differential of the complex $\mathop{\rm Tot}\nolimits^{q}\Bigl{(}\bigl{(}\mathcal{M}_{f}\big{/}\mathcal{F}^{\ell}_{\lambda}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\Omega^{0,\bullet}_{X}\Bigr{)}$.
Let us consider cohomology classes
$\alpha_{\ell}\in H^{q}\Bigl{(}X,\bigl{(}\mathcal{M}_{f}\big{/}\mathcal{F}^{\ell}_{\lambda}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}\quad(\ell=1,2,\ldots)$
such that $\pi_{\ell+1}(\alpha_{\ell+1})=\alpha_{\ell}$. We shall construct cocycles
$a_{\ell}\in H^{0}\left(X,\mathop{\rm Tot}\nolimits^{q}\Bigl{(}\bigl{(}\mathcal{M}_{f}\big{/}\mathcal{F}^{\ell}_{\lambda}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\otimes\Omega^{0,\bullet}_{X}\Bigr{)}\right)\quad(\ell=1,2,\ldots)$
[…]
Because $\pi_{\ell+1}(\alpha_{\ell+1})=\alpha_{\ell}$, there exists
[…]
such that $d(b_{\ell+1})=d(a^{\prime}_{\ell+1})$. Let $\alpha^{\prime}_{\ell+1}$ denote the cohomology class of $a^{\prime}_{\ell+1}-b_{\ell+1}$. Because $\alpha_{\ell+1}-\alpha^{\prime}_{\ell+1}$ comes from $H^{q}\Bigl{(}X,\mathop{\rm Tot}\nolimits\Bigl{(}(\mathcal{F}_{\lambda}^{\ell}\mathcal{M}_{f}\big{/}\mathcal{F}_{\lambda}^{\ell+1}\mathcal{M}_{f})\otimes\widetilde{\Omega}_{\mathcal{X}/{\mathbb{C}}}^{\bullet}\otimes\Omega_{X}^{0,\bullet}\Bigr{)}\Bigr{)}$, there exists $a_{\ell+1}$ with the desired property.
Let us consider coboundaries
$a_{\ell}\in H^{0}\Bigl{(}X,\mathop{\rm Tot}\nolimits^{q}\Bigl{(}\bigl{(}\mathcal{M}_{f}/\mathcal{F}_{\lambda}^{\ell}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{X}\otimes\Omega_{X}^{0,\bullet}\Bigr{)}\Bigr{)}\quad(\ell=1,2,\ldots)$
such that $\pi_{\ell+1}(a_{\ell+1})=a_{\ell}$. Let us construct
$b_{\ell}\in H^{0}\Bigl{(}X,\mathop{\rm Tot}\nolimits^{q-1}\Bigl{(}\bigl{(}\mathcal{M}_{f}\big{/}\mathcal{F}_{\lambda}^{\ell}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{X}\otimes\Omega_{X}^{0,\bullet}\Bigr{)}\Bigr{)}\quad(\ell=1,2,\ldots)$
such that $d(b_{\ell})=a_{\ell}$ and $\pi_{\ell+1}(b_{\ell+1})=b_{\ell}$ inductively on $\ell$. Suppose that we have already constructed $b_{\ell}$. There exists $b^{\prime}_{\ell+1}\in H^{0}\Bigl{(}X,\mathop{\rm Tot}\nolimits^{q-1}\Bigl{(}\bigl{(}\mathcal{M}_{f}\big{/}\mathcal{F}_{\lambda}^{\ell+1}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{X}\otimes\Omega_{X}^{0,\bullet}\Bigr{)}\Bigr{)}$ such that $\pi_{\ell+1}(b^{\prime}_{\ell+1})=b_{\ell}$. We obtain a cocycle
$a_{\ell+1}-d(b^{\prime}_{\ell+1})\in H^{0}\Bigl{(}X,\mathop{\rm Tot}\nolimits^{q}\Bigl{(}\bigl{(}\mathcal{F}_{\lambda}^{\ell}\mathcal{M}_{f}\big{/}\mathcal{F}_{\lambda}^{\ell+1}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{X}\otimes\Omega_{X}^{0,\bullet}\Bigr{)}\Bigr{)}.$
By Theorem 3.6, the natural morphism
$H^{q}\bigl{(}X,\mathop{\rm Gr}\nolimits^{\ell}_{\mathcal{F}_{\lambda}}\mathcal{M}\otimes\widetilde{\Omega}_{\mathcal{X}/{\mathbb{C}}}^{\bullet}\bigr{)}\longrightarrow H^{q}\bigl{(}X,(\mathcal{M}/\mathcal{F}_{\lambda}^{\ell+1}\mathcal{M})\otimes\widetilde{\Omega}_{\mathcal{X}/{\mathbb{C}}}^{\bullet}\bigr{)}$
is injective. Hence, there exists
$c_{\ell+1}\in H^{0}\Bigl{(}X,\mathop{\rm Tot}\nolimits^{q-1}\Bigl{(}\bigl{(}\mathcal{F}_{\lambda}^{\ell}\mathcal{M}_{f}\big{/}\mathcal{F}_{\lambda}^{\ell+1}\mathcal{M}_{f}\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{X}\otimes\Omega_{X}^{0,\bullet}\Bigr{)}\Bigr{)}$
such that $d(c_{\ell+1})=a_{\ell+1}-d(b^{\prime}_{\ell+1})$. By setting $b_{\ell+1}=b^{\prime}_{\ell+1}+c_{\ell+1}$, the induction can proceed.
3.2.3 Reformulation
Let $\iota_{f}:X\to X\times{\mathbb{C}}_{t}$ denote the graph embedding. We obtain the $\widetilde{\mathcal{R}}_{X\times{\mathbb{C}}_{t}}$-module $\widetilde{\mathcal{M}}=\iota_{f\dagger}\mathcal{M}$ on $X\times{\mathbb{C}}_{t}$. The following lemma is standard.
Lemma 3.13
There exists a natural isomorphism $\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\simeq\iota_{f\dagger}(\mathcal{M}_{f})$. As a result, there exists a natural quasi-isomorphism:
$\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}\simeq\iota_{f\ast}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\bigr{)}[-1].$
It induces the quasi-isomorphisms of the subcomplexes
$\mathcal{F}^{k}_{\lambda}\bigl{(}\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}\bigr{)}\simeq\iota_{f\ast}\bigl{(}\mathcal{F}^{k}_{\lambda}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\bigr{)}\bigr{)}[-1]$
and the quotient complexes:
$\mathop{\rm Gr}\nolimits^{k}_{\mathcal{F}_{\lambda}}\bigl{(}\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}\bigr{)}\simeq\iota_{f\ast}\bigl{(}\mathop{\rm Gr}\nolimits^{k}_{\mathcal{F}_{\lambda}}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\bigr{)}\bigr{)}[-1].$
[…]
  • $\mathop{\rm Gr}\nolimits^{V}_{a}=V_{a}/V_{<a}$ are flat over $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$.
  • $tV_{a}\subset V_{a-1}$, and $tV_{a}=V_{a-1}$ if $a<0$.
  • $\eth_{t}V_{a}\subset V_{a+1}$, and the induced morphisms $\eth_{t}:\mathop{\rm Gr}\nolimits^{V}_{a}\longrightarrow\mathop{\rm Gr}\nolimits^{V}_{a+1}$ are isomorphisms if a>1a>-1.
[…]
We obtain the following complex:
$V_{-1}(\widetilde{\mathcal{M}})\longrightarrow\lambda^{-1}V_{0}(\widetilde{\mathcal{M}})\otimes dt,\quad s\longmapsto(\partial_{t}+\lambda^{-1})s\,dt.$
The first term sits in the degree 0. It extends to the following double complex:
$V_{-1}(\widetilde{\mathcal{M}})\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\longrightarrow\lambda^{-1}V_{0}(\widetilde{\mathcal{M}})\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}.$ (6)
Let $\mathfrak{C}_{1}(\mathcal{M})$ denote the complex of sheaves on ${\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}$ obtained as the total complex of (6). We have the subcomplexes $\mathcal{F}^{j}_{\lambda}\mathfrak{C}_{1}(\mathcal{M})=\lambda^{j}\mathfrak{C}_{1}(\mathcal{M})$. We may naturally regard $\mathfrak{C}_{1}(\mathcal{M})$ as a subcomplex of $\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}$.
Lemma 3.14
The inclusion $\mathfrak{C}_{1}(\mathcal{M})\to\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}$ is a quasi-isomorphism. It induces quasi-isomorphisms of subcomplexes
$\mathcal{F}_{\lambda}^{k}\mathfrak{C}_{1}(\mathcal{M})\longrightarrow\mathcal{F}_{\lambda}^{k}\bigl{(}\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}\bigr{)}$
and the quotient complexes:
$\mathop{\rm Gr}\nolimits^{k}_{\mathcal{F}_{\lambda}}\mathfrak{C}_{1}(\mathcal{M})\longrightarrow\mathop{\rm Gr}\nolimits^{k}_{\mathcal{F}_{\lambda}}\bigl{(}\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}\bigr{)}.$
Proof  For any a>1a>-1, the induced morphisms
$\mathop{\rm Gr}\nolimits^{V}_{a}(\widetilde{\mathcal{M}})\to\lambda^{-1}\mathop{\rm Gr}\nolimits^{V}_{a+1}(\widetilde{\mathcal{M}})\otimes dt\quad s\longmapsto\partial_{t}(s)\otimes dt$
are isomorphisms. We obtain that the quotient of $\mathfrak{C}_{1}(\mathcal{M})\to\widetilde{\mathcal{M}}\otimes\mathcal{L}(t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}$ is acyclic, that is the first claim. The second claim is equivalent to the first claim. The third claim follows from the second claim.
By Lemma 3.14, Theorem 3.6 is reduced to the following theorem.
Theorem 3.15
The $E_{1}$-degeneration property holds for the filtered complex $\mathcal{F}_{\lambda}^{\bullet}\mathfrak{C}_{1}(\mathcal{M})$ with respect to the push-forward by the projection ${\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\to{\mathbb{C}}_{\lambda}$.
3.2.4 Refinement
Let $\pi_{012}:{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\to{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}$ and $\pi_{01}:{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\to{\mathbb{C}}_{\lambda}\times X$ denote the projections. We obtain the following complex on ${\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}$:
$\pi_{012}^{\ast}V_{-1}(\widetilde{\mathcal{M}})\longrightarrow\lambda^{-1}\pi_{012}^{\ast}V_{0}(\widetilde{\mathcal{M}})\otimes dt,\quad s\longmapsto(u\partial_{t}+\lambda^{-1})s\,dt.$ (7)
The first term sits in the degree 0. We extend it to the following double complex:
$\pi_{012}^{\ast}V_{-1}(\widetilde{\mathcal{M}})\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\longrightarrow\lambda^{-1}\pi_{012}^{\ast}V_{0}(\widetilde{\mathcal{M}})\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}.$
Let $\mathfrak{C}(\mathcal{M})$ denote the associated complex on ${\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}$. We have the subcomplexes $\mathcal{F}_{\lambda}^{j}\mathfrak{C}(\mathcal{M})=\lambda^{j}\mathfrak{C}(\mathcal{M})$.
[…]
Theorem 3.16
[…]
We obtain Theorem 3.15 from Theorem 3.16 by specializing along $u=1$.
3.3 Formal neighbourhood along $u=0$
3.3.1 Basic strictness
Let $\pi_{0}:{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\to{\mathbb{C}}_{\lambda}$ denote the projection. Let $\iota_{u}:{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times\{0\}\to{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}$ and $\iota_{01}:{\mathbb{C}}_{\lambda}\times X\times\{0\}\longrightarrow{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}$ denote the inclusion maps.
Lemma 3.17
$R^{\ell}\pi_{0\ast}\bigl{(}\iota_{u}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}$ are locally free $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$-modules of finite rank for any $\ell$.
[…]
Note that $\iota_{u}^{\ast}\mathfrak{C}(\mathcal{M})$ is quasi-isomorphic to
$\lambda^{-1}\iota_{01\ast}\Bigl{(}\bigl{(}V_{0}(\widetilde{\mathcal{M}})/V_{-1}(\widetilde{\mathcal{M}})\bigr{)}\otimes dt\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}[-1].$
For any $b<a$, there exists the following canonical splitting of $V\mathcal{R}_{X\times{\mathbb{C}}_{t}|{\mathbb{C}}_{\lambda}^{\ast}\times X\times{\mathbb{C}}_{t}}$-modules
$\bigl{(}V_{a}(\widetilde{\mathcal{M}})/V_{b}(\widetilde{\mathcal{M}})\bigr{)}_{|{\mathbb{C}}_{\lambda}^{\ast}\times X\times{\mathbb{C}}_{t}}=\bigoplus_{b<c\leq a}\mathop{\rm Gr}\nolimits^{V}_{c}(\widetilde{\mathcal{M}})_{|{\mathbb{C}}_{\lambda}^{\ast}\times X\times{\mathbb{C}}_{t}}.$ (8)
Here, the actions of $\partial_{t}t-c$ on $\mathop{\rm Gr}\nolimits^{V}_{c}(\widetilde{\mathcal{M}})_{|{\mathbb{C}}_{\lambda}^{\ast}\times X\times{\mathbb{C}}_{t}}$ are nilpotent. For any $1<b<a\leq 0$, we obtain the following exact sequence for any $\ell$:
$0\longrightarrow R^{\ell}\pi_{0\ast}\bigl{(}\mathop{\rm Gr}\nolimits^{V}_{b}(\widetilde{\mathcal{M}})\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\bigr{)}_{|{\mathbb{C}}_{\lambda}^{\ast}}\longrightarrow R^{\ell}\pi_{0\ast}\Bigl{(}\bigl{(}V_{a}(\widetilde{\mathcal{M}})/V_{<b}(\widetilde{\mathcal{M}})\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\Bigr{)}_{|{\mathbb{C}}_{\lambda}^{\ast}}\\ \longrightarrow R^{\ell}\pi_{0\ast}\Bigl{(}\bigl{(}V_{a}(\widetilde{\mathcal{M}})/V_{b}(\widetilde{\mathcal{M}})\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\Bigr{)}_{|{\mathbb{C}}_{\lambda}^{\ast}}\longrightarrow 0.$ (9)
Because $R^{k}\pi_{0\ast}\bigl{(}\mathop{\rm Gr}\nolimits^{V}_{b}(\widetilde{\mathcal{M}})\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\bigr{)}$ are locally free $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$-modules for any $k$, we obtain the vanishing of the following morphisms:
$R^{\ell}\pi_{0\ast}\Bigl{(}\bigl{(}V_{a}(\widetilde{\mathcal{M}})/V_{b}(\widetilde{\mathcal{M}})\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\Bigr{)}\longrightarrow R^{\ell+1}\pi_{0\ast}\bigl{(}\mathop{\rm Gr}\nolimits^{V}_{b}(\widetilde{\mathcal{M}})\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\bigr{)}.$
We obtain the following exact sequences:
$0\longrightarrow R^{\ell}\pi_{0\ast}\bigl{(}\mathop{\rm Gr}\nolimits^{V}_{b}(\widetilde{\mathcal{M}})\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\bigr{)}\longrightarrow R^{\ell}\pi_{0\ast}\Bigl{(}\bigl{(}V_{a}(\widetilde{\mathcal{M}})/V_{<b}(\widetilde{\mathcal{M}})\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\Bigr{)}\\ \longrightarrow R^{\ell}\pi_{0\ast}\Bigl{(}\bigl{(}V_{a}(\widetilde{\mathcal{M}})/V_{b}(\widetilde{\mathcal{M}})\bigr{)}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\Bigr{)}\longrightarrow 0.$ (10)
Then, we obtain the claim of Lemma 3.17 by an easy induction.
3.3.2 The $E_{1}$-degeneration for another filtration $\mathcal{F}_{u}$
We consider the subcomplexes $\mathcal{F}_{u}^{j}=u^{j}\mathfrak{C}(\mathcal{M})$. There exist natural isomorphisms $\mathop{\rm Gr}\nolimits_{\mathcal{F}_{u}}^{j}\mathfrak{C}(\mathcal{M})\simeq\iota_{u\ast}\iota_{u}^{\ast}\mathfrak{C}(\mathcal{M})$. There exists the following exact sequence:
$0\longrightarrow\mathop{\rm Gr}\nolimits_{\mathcal{F}_{u}}^{j}\mathfrak{C}(\mathcal{M})\longrightarrow\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j+1}\mathfrak{C}(\mathcal{M})\longrightarrow\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\longrightarrow 0.$ (11)
Proposition 3.18
We obtain the following exact sequences:
$0\longrightarrow R^{\ell}\pi_{03\ast}\mathop{\rm Gr}\nolimits_{\mathcal{F}_{u}}^{j}\mathfrak{C}(\mathcal{M})\longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j+1}\mathfrak{C}(\mathcal{M})\bigr{)}\\ \longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}\longrightarrow 0.$ (12)
Proof  Let us consider the following exact sequence:
$0\longrightarrow\mathcal{F}_{u}^{1}\mathfrak{C}(\mathcal{M})\big{/}\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\longrightarrow\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\\ \longrightarrow\mathop{\rm Gr}\nolimits_{\mathcal{F}_{u}}^{0}\mathfrak{C}(\mathcal{M})\longrightarrow 0.$
Lemma 3.19
We obtain the following exact sequences for any $\ell\in{\mathbb{Z}}_{\geq 0}$ and $j\in{\mathbb{Z}}_{\geq 1}$:
$0\longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathcal{F}_{u}^{1}\mathfrak{C}(\mathcal{M})\big{/}\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\\ \longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathop{\rm Gr}\nolimits_{\mathcal{F}_{u}}^{0}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\longrightarrow 0.$ (13)
Proof  Recall that $\mathop{\rm Gr}\nolimits^{0}_{\mathcal{F}_{u}}\mathfrak{C}(\mathcal{M})=\iota_{u\ast}\iota_{u}^{\ast}\mathfrak{C}(\mathcal{M})$ is quasi-isomorphic to
$\iota_{u\ast}\Bigl{(}\lambda^{-1}\iota_{01\ast}\Bigl{(}\pi_{012}^{\ast}\bigl{(}V_{0}(\widetilde{\mathcal{M}})/V_{-1}(\widetilde{\mathcal{M}})\bigr{)}\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\Bigr{)}[-1]\Bigr{)}.$
Let $\pi_{1}:{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\to X$ denote the projection. Let $\mathcal{C}_{2}^{\bullet}$ denote the Dolbeault resolution, i.e., the associated complex of the double complex
$\iota_{u\ast}\Bigl{(}\lambda^{-1}\iota_{01\ast}\Bigl{(}\pi_{012}^{\ast}\bigl{(}V_{0}(\widetilde{\mathcal{M}})/V_{-1}(\widetilde{\mathcal{M}})\bigr{)}\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}\Bigr{)}[-1]\Bigr{)}\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}.$
For any $N\geq 0$, there exists the following subcomplex $V_{-N-1}\mathfrak{C}(\mathcal{M})$ of $\mathfrak{C}(\mathcal{M})$:
$\pi_{012}^{\ast}V_{-N-1}(\widetilde{\mathcal{M}})\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\longrightarrow\lambda^{-1}\pi_{012}^{\ast}V_{-N}(\widetilde{\mathcal{M}})\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}.$
Let $U$ be an open subset of ${\mathbb{C}}_{\lambda}^{\ast}\times{\mathbb{C}}_{u}$. Let $\alpha$ be a section of $\mathcal{C}_{2}^{\bullet}$ on $U\times X\times{\mathbb{C}}_{t}$ such that $d\alpha=0$. Let $N>j+10$. By using the splitting (8), we can construct a section $\widetilde{\alpha}$ of $\Bigl{(}\lambda^{-1}\pi_{012}^{\ast}V_{0}(\widetilde{\mathcal{M}})\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}[-1]\Bigr{)}\otimes\pi_{1}^{\ast}\Omega^{0,\bullet}_{X}$ such that (i) $\widetilde{\alpha}$ induces $\alpha$, (ii) $\beta^{(0)}:=d\widetilde{\alpha}$ is contained in $\Bigl{(}\lambda^{-1}\pi_{012}^{\ast}V_{-N}(\widetilde{\mathcal{M}})\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}$ on $U\times X\times{\mathbb{C}}_{t}$. Note that $d\beta^{(0)}=0$. There exists a section $\gamma^{(1)}$ of $\Bigl{(}\pi_{012}^{\ast}V_{-N}(\widetilde{\mathcal{M}})\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}$ on $U\times X\times{\mathbb{C}}_{t}$ such that $\gamma^{(1)}\otimes\lambda^{-1}dt=\beta^{(0)}$. We set $\beta^{(1)}=\beta^{(0)}-d\gamma^{(1)}$ which is contained in
$u\Bigl{(}\lambda^{-1}\pi_{012}^{\ast}V_{-N+1}(\widetilde{\mathcal{M}})\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}$
on $U\times X\times{\mathbb{C}}_{t}$. We have $d\beta^{(1)}=0$. Inductively, for $m=1,\ldots,j+1$, we can construct sections $\beta^{(m)}$ of
$u^{m}\Bigl{(}\lambda^{-1}\pi_{012}^{\ast}V_{-N+m}(\widetilde{\mathcal{M}})\otimes dt\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}$
and sections $\gamma^{(m)}$ of $u^{m-1}\Bigl{(}\pi_{012}^{\ast}V_{-N+m-1}(\widetilde{\mathcal{M}})\otimes\pi_{01}^{\ast}\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\Bigr{)}\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}$ such that $\beta^{(m)}=\beta^{(m-1)}-d\gamma^{(m)}$ and $\gamma^{(m)}\otimes\lambda dt=\beta^{(m-1)}$. Then, $\hat{\alpha}=\widetilde{\alpha}-\sum_{m=1}^{j+1}\gamma^{(m)}$ is a section of $\mathfrak{C}(\mathcal{M})\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}$ such that (i) $\hat{\alpha}$ induces $\alpha$, (ii) $d\hat{\alpha}$ is a section of $\mathcal{F}_{u}^{j+1}\mathfrak{C}(\mathcal{M})\otimes\pi_{1}^{\ast}\Omega_{X}^{0,\bullet}[1]$. Then, we obtain the claim of Lemma 3.19.
Lemma 3.20
We obtain the following exact sequences for any $\ell,j\in{\mathbb{Z}}_{\geq 0}$.
$0\longrightarrow R^{\ell}\pi_{03\ast}\mathop{\rm Gr}\nolimits^{j}_{\mathcal{F}_{u}}\mathfrak{C}(\mathcal{M})_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j+1}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\\ \longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\longrightarrow 0.$ (14)
Proof  It is enough to prove that
$R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j+1}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}$ (15)
are epimorphisms for any $\ell$ and $j$. We use an induction on $j$. Lemma 3.19 implies the claim in the case $j=1$. By the hypothesis of the induction, we may assume that
$R^{\ell}\pi_{03\ast}\bigl{(}\mathcal{F}_{u}^{1}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j+1}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}\longrightarrow R^{\ell}\pi_{03\ast}\bigl{(}\mathcal{F}_{u}^{1}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}_{|{\mathbb{C}}^{\ast}_{\lambda}\times{\mathbb{C}}_{u}}$
are epimorphisms. Then, by Lemma 3.19, we obtain that (15) are also epimorphisms.
We obtain the exactness of (12) from the exactness of (14) and the strictness of $R^{\ell}\pi_{03\ast}\mathop{\rm Gr}\nolimits_{\mathcal{F}_{u}}^{j}\mathfrak{C}(\mathcal{M})$ for any $\ell$ and $j$ as in the proof of Lemma 3.17. Thus, we obtain Proposition 3.18.
The supports of $R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}$ are contained in ${\mathbb{C}}_{\lambda}\times\{0\}\subset{\mathbb{C}}_{\lambda}\times{\mathbb{C}}_{u}$. We may naturally regard them as $\mathcal{O}_{{\mathbb{C}}_{\lambda}}[u]/u^{j}$-modules.
Corollary 3.21
$R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}$ are locally free $\mathcal{O}_{{\mathbb{C}}_{\lambda}}[u]/u^{j}$-modules.
3.4 Coherence of the specialization along $\lambda=0$
Let $\widetilde{\iota}_{\lambda}:\{0\}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\longrightarrow{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}$ denote the inclusion. Let $\pi_{3}:X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\to{\mathbb{C}}_{u}$ denote the projection.
Proposition 3.22
$R^{\ell}\pi_{3\ast}(\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M}))$ are coherent $\mathcal{O}_{{\mathbb{C}}_{u}}$-modules.
Proof  Let $\pi_{12}:X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\to X\times{\mathbb{C}}_{t}$ denote the projection. Note that $\widetilde{\iota}_{\lambda}^{\ast}\pi_{12}^{\ast}\mathcal{R}_{X\times{\mathbb{C}}_{t}}$ is isomorphic to the algebra of the symmetric product of $\pi_{12}^{\ast}\Theta_{X\times{\mathbb{C}}_{t}}$, where $\Theta_{X\times{\mathbb{C}}_{t}}$ denote the tangent sheaf of $X\times{\mathbb{C}}_{t}$. Because $\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})$ is a complex of coherent $\widetilde{\iota}_{\lambda}^{\ast}\mathcal{R}_{X\times{\mathbb{C}}_{t}}$-modules, it induces a complex of coherent $\mathcal{O}_{T^{\ast}(X\times{\mathbb{C}}_{t})\times{\mathbb{C}}_{u}}$-modules $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}^{\sim}$. Let $Z$ denote the cohomological support of $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}^{\sim}$.
Lemma 3.23
Let $0_{X\times{\mathbb{C}}_{t}}:X\times{\mathbb{C}}_{t}\to T^{\ast}(X\times{\mathbb{C}}_{t})$ denote the 0-section. Then, $Z\subset\bigl{(}0_{X\times{\mathbb{C}}_{t}}\circ\iota_{f}\bigr{)}(\mathop{\rm Cr}\nolimits(f))\times{\mathbb{C}}_{u}$.
Proof  Let $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}_{u}$ denote the pull back of $\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})$ by the inclusion $X\times{\mathbb{C}}_{t}\times\{u\}\to X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}$. Let $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}_{u}^{\sim}$ denote the induced complex of coherent $\mathcal{O}_{T^{\ast}(X\times{\mathbb{C}}_{t})}$-modules. It equals the pull back of $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}^{\sim}$ by the inclusion $T^{\ast}(X\times{\mathbb{C}}_{t})\times\{u\}\to T^{\ast}(X\times{\mathbb{C}}_{t})\times{\mathbb{C}}_{u}$. Let $Z_{u}$ denote the cohomological support of $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}_{u}^{\sim}$. Because $Z\cap\bigl{(}T^{\ast}(X\times{\mathbb{C}}_{t})\times\{u\}\bigr{)}=Z_{u}$, it is enough to prove that $Z_{u}\subset(0_{X\times{\mathbb{C}}_{t}}\circ\iota_{f})(\mathop{\rm Cr}\nolimits(f))$.
Let $\iota_{0,\lambda}:\{0\}\times X\times{\mathbb{C}}_{t}\to{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}$ denote the inclusion. By Lemma 3.14, if $u\neq 0$, $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}_{u}$ is quasi-isomorphic to $\iota_{0,\lambda}^{\ast}\Bigl{(}\widetilde{\mathcal{M}}\otimes\mathcal{L}(u^{-1}t)\otimes\widetilde{\Omega}^{\bullet}_{{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}/{\mathbb{C}}_{\lambda}}\Bigr{)}$. It is quasi-isomorphic to $\iota_{f\ast}\iota_{\lambda}^{\ast}\Bigl{(}\mathcal{M}_{u^{-1}f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}_{\lambda}}[-1]\Bigr{)}$. Hence, by Lemma 3.1, the cohomological support of $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}^{\sim}_{u}$ is contained in $(0_{X\times{\mathbb{C}}_{t}}\circ\iota_{f})(\mathop{\rm Cr}\nolimits(f))$.
Let us consider the case $u=0$. There exists the following quasi-isomorphism
$\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}_{0}\simeq\iota_{\lambda^{\ast}}\bigl{(}V_{0}(\widetilde{\mathcal{M}})/V_{-1}(\widetilde{\mathcal{M}})\bigr{)}\otimes dt\otimes\Omega^{\bullet}_{X}[-1].$
Because $df(X)\cap\mathop{\rm Ch}\nolimits(\Xi_{\mathop{\rm DR}\nolimits}(\mathcal{M}))\subset 0_{X}(\mathop{\rm Cr}\nolimits(f))$, the support of the $\mathop{\rm Sym}\nolimits\Theta_{X}$-module $)\iota_{\lambda\ast}(V_{0}(\widetilde{\mathcal{M}})/V_{-1}(\widetilde{\mathcal{M}}$)) is contained in $\iota_{f}(\mathop{\rm Cr}\nolimits(f))$ as in the case of Lemma 3.2. By a similar argument to the proof of Lemma 3.1, we obtain that the cohomological support of $\bigl{(}\widetilde{\iota}_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})\bigr{)}_{0}$ is contained in $(0_{X\times{\mathbb{C}}_{t}}\circ\iota_{f})(\mathop{\rm Cr}\nolimits(f))$.
Because $Z$ is proper over ${\mathbb{C}}_{u}$, we obtain the claim of Proposition 3.22.
Note that $\mathop{\rm Gr}\nolimits^{j}_{\mathcal{F}_{\lambda}}\mathfrak{C}(\mathcal{M})$ are isomorphic to $\iota_{\lambda\ast}\iota_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})$ for any $j\in{\mathbb{Z}}_{\geq 0}$. The supports of $R^{\ell}\pi_{03\ast}(\mathfrak{C}(\mathcal{M})/\mathcal{F}^{j}_{\lambda}\mathfrak{C}(\mathcal{M}))$ are contained in $\{0\}\times{\mathbb{C}}_{u}$. We may naturally regard $R^{\ell}\pi_{03\ast}(\mathfrak{C}(\mathcal{M})/\mathcal{F}^{j}_{\lambda}\mathfrak{C}(\mathcal{M}))$ as $\mathcal{O}_{{\mathbb{C}}_{u}}$-modules.
Corollary 3.24
$R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{\lambda}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}$ are coherent $\mathcal{O}_{{\mathbb{C}}_{u}}$-modules for any $\ell$ and $j$.
3.5 Proof of Theorem 3.16
There exist the following exact sequences:
$0\longrightarrow\lambda^{j}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Big{/}\lambda^{j+1}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\\ \longrightarrow\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Big{/}\lambda^{j+1}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\\ \longrightarrow\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Big{/}\lambda^{j}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\longrightarrow 0.$ (16)
We obtain the following morphisms of the stalks at $(0,0)\in{\mathbb{C}}_{\lambda}\times{\mathbb{C}}_{u}$:
$R^{\ell}\pi_{03\ast}\Bigl{(}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Big{/}\lambda^{j}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Bigr{)}_{(0,0)}\longrightarrow\\ R^{\ell+1}\pi_{03\ast}\Bigl{(}\lambda^{j}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Big{/}\lambda^{j+1}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Bigr{)}_{(0,0)}.$ (17)
Lemma 3.25
The morphisms (17) are 0.
Proof  Because the multiplication of $\lambda^{j}$ on $R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}$ is a monomorphism by Corollary 3.21, the induced morphism
$R^{\ell}\pi_{03\ast}\Bigl{(}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\big{/}\lambda^{j}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Bigr{)}_{(0,0)}\longrightarrow R^{\ell+1}\pi_{03\ast}\Bigl{(}\lambda^{j}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})\bigr{)}\Bigr{)}_{(0,0)}$
is 0. Hence, the morphism (17) is 0.
There exist the following exact sequences:
$0\longrightarrow\mathop{\rm Gr}\nolimits_{\mathcal{F}_{\lambda}}^{j}\mathfrak{C}(\mathcal{M})\longrightarrow\mathfrak{C}(\mathcal{M})/\mathcal{F}^{j+1}_{\lambda}\mathfrak{C}(\mathcal{M})\longrightarrow\mathfrak{C}(\mathcal{M})/\mathcal{F}^{j}_{\lambda}\mathfrak{C}(\mathcal{M})\longrightarrow 0.$
We obtain the following induced morphisms of the stalks:
$R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}^{j}_{\lambda}\mathfrak{C}(\mathcal{M})\bigr{)}_{(0,0)}\longrightarrow R^{\ell+1}\pi_{03\ast}\bigl{(}\mathop{\rm Gr}\nolimits_{\mathcal{F}_{\lambda}}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}_{(0,0)}.$ (18)
Lemma 3.26
The morphisms (18) are 0.
Proof  There exist the following commutative diagrams of stalks at $(\lambda,u)=(0,0)$ for any $k\geq 0$:
[…]
The lower horizontal arrow is 0 by Lemma 3.25. By the construction, $\mathfrak{C}(\mathcal{M})$ is flat over $\mathcal{O}_{{\mathbb{C}}_{\lambda}\times{\mathbb{C}}_{u}}$, and hence $\lambda^{p}\mathfrak{C}(\mathcal{M})\cap u^{q}\mathfrak{C}(\mathcal{M})=\lambda^{p}u^{q}\mathfrak{C}(\mathcal{M})$ for any $p,q\in{\mathbb{Z}}_{\geq 0}$. The right vertical arrows are identified as follows:
[…]
Under the identification, the image of $b^{\ell}$ is contained in the image of the morphism
$R^{\ell+1}\pi_{03\ast}\iota_{\lambda\ast}\iota_{\lambda}^{\ast}\mathcal{F}_{u}^{k}\mathfrak{C}(\mathcal{M})_{(0,0)}\longrightarrow R^{\ell+1}\pi_{03\ast}\iota_{\lambda\ast}\iota_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})_{(0,0)}$
for any $k\geq 0$. It equals the image of the morphism
$u^{k}:R^{\ell+1}\pi_{03\ast}\iota_{\lambda\ast}\iota_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})_{(0,0)}\longrightarrow R^{\ell+1}\pi_{03\ast}\iota_{\lambda\ast}\iota_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})_{(0,0)}$
for any $k\geq 0$. Because $R^{\ell+1}\pi_{03\ast}\iota_{\lambda\ast}\iota_{\lambda}^{\ast}\mathfrak{C}(\mathcal{M})$ is $\mathcal{O}_{{\mathbb{C}}_{\lambda}}$-coherent by Proposition 3.22, we obtain $b^{\ell}=0$.
For any $j\geq 0$, there exists $\epsilon>0$ such that the following morphism is an epimorphism on $\{|u|<\epsilon\}$:
$R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{\lambda}^{j+1}\mathfrak{C}(\mathcal{M})\bigr{)}\longrightarrow\\ R^{\ell}\pi_{03\ast}\bigl{(}\mathfrak{C}(\mathcal{M})/\mathcal{F}_{\lambda}^{j}\mathfrak{C}(\mathcal{M})\bigr{)}.$
[…]